# Moran Mizrahi1 Chen Shani2 Gabriel Stanovsky1 Dan Jurafsky2 Dafna Shahaf1 1

### v1

@@ Abstract
Large Language Models (LLMs) excel at countless tasks, yet struggle with creativity. In this paper, we introduce a novel approach that couples LLMs with structured representations and cognitively inspired manipulations to generate more creative and diverse ideas. Our notion of creativity goes beyond superficial token-level variations; rather, we explicitly recombine structured representations of existing ideas, allowing our algorithm to effectively explore the more abstract landscape of ideas. We demonstrate our approach in the culinary domain with $\mathcal{D}$ishCover, a model that generates creative recipes. Experiments comparing our model’s results to those of GPT-4o show greater diversity. Domain expert evaluations reveal that our outputs, which are mostly coherent and feasible culinary creations, significantly surpass GPT-4o in terms of novelty, thus outperforming it in creative generation. We hope our work inspires further research into structured creativity in AI.
[Uncaptioned image] Cooking Up Creativity: A Cognitively-Inspired Approach for Enhancing LLM Creativity through Structured Representations
Moran Mizrahi1 Chen Shani2 Gabriel Stanovsky1 Dan Jurafsky2 Dafna Shahaf1 1Hebrew University of Jerusalem  2Stanford University {moranmiz, dshahaf, gabriel.stanovsky}@cs.huji.ac.il  {cshani, jurafsky}@stanford.edu
@@ 1 Introduction
Large Language Models (LLMs) excel at generating fluent, coherent text and performing tasks that draw on extensive world knowledge. However, they often struggle to generate truly creative ideas Franceschelli and Musolesi (2024); Chakrabarty et al. (2024); Tian et al. (2024b); Zhao et al. (2024).
@@
In creativity research, creative outputs are typically defined as those that are both novel (unexpected and original) and valuable (useful, relevant, or effective) Mumford (2003); Boden (2004). However, due to their reliance on vast repositories of existing data, LLMs inherently follow learned patterns, making them prone to producing repetitive or predictable outputs that lack genuine novelty. Ironically, attempts to explicitly instruct LLMs to “think more creatively” often leads them to generate invalid or hallucinated solutions that could mislead uninformed users Wang et al. (2024a); Jiang et al. (2024). Together, these limitations make creative generation a persistent challenge for LLMs.
@@
The temperature parameter of LLMs controls the amount of randomness, and is often claimed to be the creativity parameter. However, creativity encompasses much more than mere randomness; a recent study Peeperkorn et al. (2024) found that while higher temperatures weakly correlate with increased novelty, their actual influence on overall creativity remains subtle and limited.
@@
Much recent work has shown that combining structured knowledge (such as knowledge graphs) with LLMs can significantly improve their performance Pan et al. (2024); Wang et al. (2024b); Feng et al. (2023); Sun et al. (2023). In particular, a line of works uses an LLM to parse text into structured representation, manipulates this representation, and (optionally) applies the LLM again to translate the result into natural language, especially in inference and reasoning tasks Yang et al. (2023); Besta et al. (2024); Zelikman et al. (2023); Zhang et al. (2025).
@@
In this work, we show that surprisingly, incorporating structure can also improve the LLM’s creativity and diversity. We stress that we do not mean creativity and diversity on the lexical (token) level; Rather, we want the model to be creative on a more abstract level, within the realm of concepts (or the “landscape of ideas”, so to speak).
@@
Our approach is illustrated in Figure 1. Similar to parsing-based approaches Tian et al. (2024a); Zhao et al. (2023); Li et al. (2024), we start by deriving structured representations from textual inputs. Next, we manipulate these structured representations to reach creative parts of the idea space. We draw inspiration from the human creative process, and focus on recombination – a fundamental principle in creativity research, which posits that novel ideas often emerge by merging existing concepts in unexpected ways Guilford (1967); Utterback (1996); Ahuja and Morris Lampert (2001). For example, combining pizza preparation method with the flavors of alfajores cookies might yield a brand-new “alfajores pizza”, and combining a sofa with a bookshelf might result in multifunctional furniture. We sample ideas from the space of recombinations, and evaluate them based on novelty and value Finke et al. (1996); Sawyer and Henriksen (2024).
@@
We demonstrate our paradigm in the culinary domain, introducing $\mathcal{D}$ishCover, a model for creative generation of recipes. Figure 2 presents examples of recombinations generated by it.
Beyond the immediate scope of cooking, we believe this paradigm holds promise for extending creative and diverse generation to a wide range of domains, from product design, narrative construction, and scientific discovery to artistic composition. In summary, our contributions are:
@@
  • We introduce a novel paradigm to enhance LLM creativity by extracting structured representations from natural language and manipulating them (inspired by cognitive processes), going beyond mere token-level variability.
@@ 2 Background: Human Creativity
Recombination of Ideas. We base our work on a prominent idea-generation method: recombination, where elements from existing ideas are merged to create novel concepts Guilford (1967); Koestler (1964). Our model strategically recombines elements from pairs of existing seed ideas to spark unexpected connections.
@@
Measuring Novelty and Value Automatically.
Novelty can be assessed by identifying how uncommon an idea is within a dataset Heinen and Johnson (2018); Kenett (2019); Doboli et al. (2020). Evaluating value, however, is highly domain-dependent, often considered the “holy grail” of computational creativity Boden (2004); Ritchie (2007); Jordanous (2012). Thus, we consider value assessment as a domain-specific task.
@@ 4 Model
In this section, we introduce $\mathcal{D}$ishCover, our model for automatically generating innovative recipes.1 Figure 1 illustrates our methodology. The input consists of two seed inspirations (idea descriptions $d_{a}$, $d_{b}$) [^2] Note that more inspirations can be used if desired.. Each idea description is mapped to a set of specific recipes (instances of the idea, for example different lasagna recipes; § 4.1).
These recipes are first translated into tree representations with the help of an LLM (step (I) in Figure 1, § 4.2). To generate new ideas, we recombine these trees to produce new candidate ideas with a minimal edit-distance algorithm (step (II), § 4.3). Then, utilizing the Novelty & Value principle, we refine the candidate ideas to assess their value and rank them based on their novelty scores (step (III), § 4.4). Finally, the highest-ranked trees are translated back into natural language recipes and refined using an LLM (step (IV), § 4.5). We now provide more details about each step.
@@ 4.1 Sampling Seed Ideas (Step I)
We selected the 100 most popular dishes (e.g., chicken salad, cheesecake) that span different categories (e.g., appetizers, desserts, main courses) from the Recipe1M+ dataset Marın et al. (2021). The average number of appearances in the dataset is 2,576.33 recipes per dish.
To keep the financial costs of using an LLM manageable in the next stage of our model, we sampled 30 recipes per dish, resulting in a total of 3K recipe samples. To ensure both diversity and representativeness, we selected 15 recipes at random to capture the typical version of each dish, and 15 more to maximize diversity. Implementation of diversification can depend on the representation and the domain; in our case, we use the GMM algorithm Ravi et al. (1994) over recipe embeddings from a Sentence-BERT model fine-tuned on recipe data (see details in Appendix A).
@@ 4.2 Text to Tree (Step II)
Cooking recipes, like science experiments, assembly manuals, and game instructions, are procedural texts. These texts typically consist of a sequence of steps accompanied by the objects needed to perform them. A common way to represent procedural texts is with a tree Jermsurawong and Habash (2015); Maeta et al. (2015), where leaf nodes correspond to needed objects (ingredients, in our case), and the internal nodes represent the actions performed on them. Figure 3 shows simple lasagna and chocolate pie recipes in this tree format.
To parse recipe text into tree, we prompted GPT-4o, leveraging its extensive world knowledge and code generation capabilities with a chain-of-thought approach. See Appendix BC for full details and corresponding prompts. The total cost of generating tree representations for 3K recipes was approximately \$40.
An initial pass revealed that 1,347 (44.9%) of the resulting trees were invalid due to issues such as orphan nodes, multiple outgoing edges from a single node, or incorrect edge directions. To address this, we implemented a correction step in which we removed problematic edges and instructed the model to reconsider them. This improved validity to 95% (2,850 trees). We then evaluated the final trees using 50 random recipes. A gold-standard tree was created for each recipe, and we automatically compared the predicted trees with the gold trees in terms of node and edge matching. For nodes, we achieved a Precision of 0.985, Recall of 0.956, and F1-Score of 0.969. For edges, we obtained a Precision of 0.951, Recall of 0.909, and F1-Score of 0.93. Overall, these results demonstrate the effectiveness of our approach in translating procedural texts into a structured tree representation, making them suitable for further manipulation and analysis.
@@ 4.4 Evaluate Ideas (Step IV)
Now that we have a set of candidate innovations created through reformulation, we evaluate each generated recipe. As noted in Section 3, the evaluation function $E$ should take into account novelty and value. Specifically, we chose to view value as a constraint, and novelty as the optimization objective; i.e., we wish to rank by novelty all candidates that pass a value threshold (i.e., make sense).
In the culinary domain, a recipe is often considered creative if it introduces unexpected ingredient or technique combinations (novelty) while still producing a delicious, coherent dish (value, utility). In the following, we operationalize these criteria. Our implementation focuses on recipes, but we believe the principle can be generalized to other domains.
@@ 4.4.1 Value Constraint
To assess value (taste) of a recipe, we follow Varshney et al. (2019), and check if its ingredients pair well with each other. We use flavorDB dataset Garg et al. (2018), which catalogs taste molecules for raw ingredients. According to this work, two ingredients pair well if they share a larger proportion of taste molecules; we compute a pairing score for each pair of raw ingredients. Since this dataset does not cover processed ingredients, we also use FoodData Central Fukagawa et al. (2022) to infer their raw components. We define the pairing score between two composite ingredients as the lowest score among their constituent raw-ingredient pairs. After exploration, we chose 0.3 as our value threshold. If the score falls below this threshold, we consider the pairing problematic. We iteratively remove the ingredient with the highest number of low-scoring pairings until no further collisions remain (alternatively, one might choose to remove these candidates altogether). See details in Appendix E.
@@ 4.4.2 Ranking by Novelty
Our intuition is that novel recipes include ingredients and actions that do not often appear together. Thus, we formulate a measure of surprise, inspired by the inverse document frequency (idf) concept in tf-idf Ramos et al. (2003).
Given recipe tree $T$ with nodes $E_{T}$ (ingredients and actions), let $N_{e}$ be the number of recipes in the repository that include element $e$. For each element $e^{\prime}\in E_{T}\setminus e$ , $df_{e}(e^{\prime})$ denotes the number of recipes containing both $e,e^{\prime}$. We then define [^3] To prevent score inflation due to typos or extremely rare ingredients, we exclude any ingredient that appears exceptionally infrequently in the entire dataset.:
$idf_{e}(e\textquoteright)=log\Bigl{(}\frac{N_{e}}{df_{e}(e^{\prime})}\Bigr{)}$
A higher $idf_{e}(e\textquoteright)$ ’ ) score indicates that $e^{\prime}$ is more unique relative to recipes containing $e$.
To compute novelty of an element in $T$, we compute all $idf_{e}(e\textquoteright)$ ’ ) scores and sum the top ten. To compute the novelty of $T$ itself, we sum the top-10 element novelties. For example, in Figure 2, the chicken and mango delight recipe was ranked as novel due to ingredient combinations like mango, crescent roll, nuts and broccoli, and actions such as steaming, blanching and unrolling.
@@ 4.5 Tree to Text (Step V)
After identifying the highest-ranked trees, we convert them back into natural-language recipes using an LLM. Similar to Section 4.2, we employ a chain-of-thought approach, instructing GPT-4o to translate the tree (encoded in DOT format) into a structured recipe, including a title, ingredient list, and step-by-step instructions. We then prompt the model to refine and correct the text for coherence, quantity adjustments, ingredient consistency, and overall readability (see Appendix F for full prompts). This approach required a total cost of approximately \$42 to process 1000 recipe trees.
@@ 6 Evaluation
To answer this, we examine two key aspects. First, how the outputs compare in terms of diversity. Specifically, we investigate whether our approach mitigates the well-documented issue of repetitiveness in LLMs, which is a critical challenge in creative generation. More importantly, we assess how our outputs compare those of GPT-4o in terms of creativity. Creativity requires outputs that are both valuable (make sense) and novel (unexpected). We compare our model’s outputs with those of GPT-4o [^4] Version: gpt-4o-2024-08-06 (latest stable version of the model at the time of writing this paper). on two tasks:
Experiment 1. We evaluated GPT-4o and $\mathcal{D}$ishCover on their ability to generate creative recipes combining a given dish pair. The evaluation included randomly selected 10 dish pairs. For each pair, we selected 5 recipes generated by GPT-4o and compared them to the top 5 recipes generated by $\mathcal{D}$ishCover, resulting in a total of 50 recipes per model.
Experiment 2. To broaden the scope of our analysis, we wish to evaluate the most creative recipes GPT-4o and $\mathcal{D}$ishCover could generate in general, without limiting them to a given input pair. We used 100 recipes from each model. We instructed GPT-4o to generate 100 different creative recipes in a single chat session. For $\mathcal{D}$ishCover, we selected 100 recipes from our 5K-recipe dataset, which includes a novelty score for each generated recipe. To ensure that the dishes do not come from a very small number of inputs, we employed simulated annealing Bertsimas and Tsitsiklis (1993) to maximize recipe novelty while enforcing a constraint on the maximum number of appearances per dish.

### v2

# Structured Recombination

@@ Abstract
Large Language Models (LLMs) excel at many tasks, yet they struggle to produce truly creative, diverse ideas. In this paper, we introduce a novel approach that enhances LLM creativity. We apply LLMs for translating between natural language and structured representations, and perform the core creative leap via cognitively inspired manipulations on these representations. Our notion of creativity goes beyond superficial token-level variations; rather, we recombine structured representations of existing ideas, enabling our system to effectively explore a more abstract landscape of ideas.
We demonstrate our approach in the culinary domain with $\mathcal{D}$ishCover, a model that generates creative recipes. Experiments and domain-expert evaluations reveal that our outputs, which are mostly coherent and feasible, significantly surpass GPT-4o in terms of novelty and diversity, thus outperforming it in creative generation. We hope our work inspires further research into structured creativity in AI.
@@ 1 Introduction
Large Language Models (LLMs) excel at generating fluent, coherent text and performing tasks that draw on extensive world knowledge. However, they often struggle to generate truly creative ideas  Franceschelli and Musolesi (2024); Chakrabarty et al. (2024a); Tian et al. (2024b); Zhao et al. (2025).
@@
In creativity research, creative outputs are typically defined as those that are both novel (unexpected and original) and valuable (useful, relevant, or effective) Mumford (2003); Boden (2004). However, due to LLM relying on vast repositories of existing data, they inherently follow learned patterns, making them prone to producing predictable, repetitive outputs that lack genuine novelty. Ironically, attempts to explicitly instruct LLMs to “think more creatively” often lead them to generate invalid or hallucinated (i.e., invaluable) solutions that could mislead uninformed users Wang et al. (2024a); Jiang et al. (2024). Together, these limitations make creative generation a persistent challenge for LLMs.
@@
The temperature parameter of LLMs controls the amount of randomness, and is often claimed to be the creativity parameter, i.e., the implicit way to enhance creativity in LLM. However, creativity encompasses much more than mere randomness; a recent study Peeperkorn et al. (2024) found that while higher temperatures weakly correlate with increased novelty, their actual influence on overall creativity remains subtle and limited.
@@
Much recent work has shown that combining LLMs with structured knowledge (e.g., knowledge graphs) can significantly improve their performance, especially in inference and reasoning tasks Feng et al. (2023); Sun et al. (2023); Pan et al. (2024); Wang et al. (2024b). Several works use LLMs to parse text into structured representations, manipulate these representations, and (optionally) apply the LLM again to translate the result into text Yang et al. (2023); Zelikman et al. (2023); Besta et al. (2024); Zhang et al. (2025).
@@
In this work, we show that surprisingly, incorporating structure can also improve LLMs’ creativity and diversity. We stress that we do not mean creativity and diversity on the lexical (token) level; rather, we want the model to be creative on a more abstract level, within the realm of concepts (or the “landscape of ideas”, so to speak).
@@
Our approach is illustrated in Figure 1. Similar to parsing-based approaches, we start by deriving structured representations from textual inputs using an LLM. We then manipulate these structured representations externally, thereby generating new ideas while systematically exploring creative regions of the idea space. Inspired by the human creative process, we focus on recombination – a fundamental principle in creativity research, which posits that novel ideas often emerge by merging existing concepts in unexpected ways Guilford (1967); Utterback (1996); Ahuja and Morris Lampert (2001). For example, combining pizza preparation methods with the flavors of alfajores cookies might yield a brand new “alfajores pizza”; combining a sofa with a bookshelf might result in multifunctional furniture. To recombine structured representations of existing ideas, we employ an edit-distance algorithm, and focus on representations midway through its transformation steps. We sample from the space of recombinations, evaluating candidates for novelty and value Finke et al. (1996); Sawyer and Henriksen (2024). Those deemed most promising are then translated into natural language.
@@
We demonstrate our paradigm in the culinary domain, introducing $\mathcal{D}$ishCover, a model for creative generation of recipes. Figure 2 presents examples of recombinations generated by it. Beyond the scope of cooking, we believe this paradigm holds promise for extending creative and diverse generation to domains where suitable structured representations and value criteria can be defined (e.g., procedural texts, drug design, music generation; see Section 8). Our contributions are:
@@
  • We introduce a novel paradigm to enhance LLM creativity by extracting structured representations, applying cognitive inspired manipulations, and decoding the results back into natural language, thus going beyond superficial token-level variation.
  • We propose a new recombination operator based on edit distance, which enables controlled blending of structured ideas by partially transforming one representation into another.
@@ 2 Background: Human Creativity
Recombination of Ideas. We base our work on a prominent idea-generation method: recombination, where elements from existing ideas are merged to create novel concepts Koestler (1964); Guilford (1967). Our model strategically recombines elements from pairs of existing seed ideas to spark unexpected connections.
@@
Measuring Novelty and Value Automatically.. Novelty can be assessed by identifying how uncommon an idea is within a dataset Heinen and Johnson (2018); Kenett (2019); Doboli et al. (2020). Evaluating value, however, is highly domain-dependent, often considered the “holy grail” of computational creativity Boden (2004); Ritchie (2007); Jordanous (2012). Thus, we consider value assessment as a domain-specific task.
@@ 4 Model
In this section, we introduce $\mathcal{D}$ishCover, our model for automatically generating innovative recipes.1 Figure 1 illustrates our methodology. The input consists of two seed inspirations (idea descriptions $d_{a}$, $d_{b}$). [^2] Note that more inspirations can be used if desired. Each idea description is mapped to a set of specific recipes (instances of the idea, for example different lasagna recipes; § 4.1).
These recipes are first translated into tree representations using an LLM (step (I) in Figure 1, § 4.2). We recombine these trees to produce new candidate ideas with a minimal edit-distance algorithm (step (II), § 4.3). Then, we refine the candidate ideas to assess their value and rank them based on their novelty scores (step (III), § 4.4). Finally, the highest-ranked trees are translated back into natural language recipes using an LLM, which leverages its commonsense and world knowledge to fill in missing details and produce coherent recipes (step (IV), § 4.5). We now provide more details about each step.
@@ 4.1 Sampling Seed Ideas (Step I)
We selected the 100 most popular dishes (e.g., chicken salad, cheesecake) that span different categories (e.g., appetizers, desserts, main courses) from the Recipe1M+ dataset Marın et al. (2021). On average, each selected dish is associated with 2,576.33 recipes in the dataset.
To keep the financial costs of using an LLM manageable in the next stage of our model, we sampled 30 recipes per dish, resulting in a total of 3K recipe samples. To ensure both diversity and representativeness, we selected 15 recipes at random to capture the typical version of each dish, and 15 more to maximize diversity. Implementation of diversification can depend on the representation and the domain; in our case, we use the GMM algorithm over recipe embeddings obtained from a Sentence-BERT model fine-tuned on recipe data (see details in Appendix A), which identifies the dish’s embedding centroid and iteratively selects recipes farthest from both the centroid and previously chosen samples Ravi et al. (1994).
@@ 4.2 Text to Tree (Step II)
Cooking recipes, like experiments, assembly manuals, and game instructions, are procedural texts. These texts typically consist of a sequence of steps accompanied by the objects needed to perform them. A common way to represent procedural texts is as a tree Jermsurawong and Habash (2015); Maeta et al. (2015), where leaf nodes correspond to needed objects (in our case, ingredients), and internal nodes represent the actions performed on them. Figure 3 shows simple lasagna and chocolate pie recipes represented as trees.
To parse recipe text into tree, we prompted GPT-4o with a chain-of-thought approach. See Appendix BC for full details and corresponding prompts. The total cost of generating tree representations for 3K recipes was approximately \$40.
An initial pass revealed that 1,347 (44.9%) of the resulting trees were invalid due to issues such as orphan nodes, multiple outgoing edges from a single node, or incorrect edge directions. To address this, we implemented a correction step where we removed problematic edges and instructed the model to reconsider them. This improved validity to 95% (2,850 trees). We then evaluated the final trees using 50 random recipes. A gold-standard tree was created for each recipe, and we automatically compared the predicted trees with the gold trees in terms of node and edge matching. For nodes, we achieved 0.985 precision, 0.956 recall, and F1-Score of 0.969. For edges, we obtained 0.951 precision, 0.909 recall, and F1-Score of 0.93. Overall, these results demonstrate the effectiveness of our approach in translating procedural texts into a structured tree representation, making them suitable for further manipulation and analysis.
@@ 4.4 Evaluate Ideas (Step IV)
Now that we have a set of candidate innovations created through recombination, we evaluate each generated recipe. As noted in Section 3, the evaluation function $E$ should take into account novelty and value. Specifically, we chose to view value as a constraint, and novelty as the optimization objective; i.e., we wish to rank by novelty all candidates that pass a value threshold (i.e., make sense).
In the culinary domain, a recipe is considered creative if it introduces unexpected ingredient or technique combinations (novelty) while resulting in a delicious, coherent dish (value, utility). We next operationalize these criteria. Our implementation is specific to recipes, but we believe the principle can be generalized to other domains.
@@ 4.4.1 Value Constraint
To assess value (taste) of a recipe, we follow Varshney et al. (2019), and check if its ingredients pair well with each other. We use flavorDB dataset Garg et al. (2018), which catalogs taste molecules for raw ingredients. According to this work, two ingredients pair well if they share a larger proportion of taste molecules; we compute a Jaccard-based pairing score for each pair of raw ingredients. Since this dataset does not cover processed ingredients, we also use FoodData Central Fukagawa et al. (2022) to infer their raw components. We define the pairing score between two composite ingredients as the lowest score among their constituent raw-ingredient pairs. After exploration, we chose 0.3 as our value threshold. If the score falls below this threshold, we consider the pairing problematic. We iteratively remove the ingredient with the highest number of low-scoring pairings until no further collisions remain (alternatively, one might choose to remove these candidates altogether). See details in Appendix E.
@@ 4.4.2 Ranking by Novelty
Our intuition is that novel recipes include ingredients and actions that do not often appear together. Thus, we formulate a measure of surprise, inspired by the inverse document frequency (idf) concept in tf-idf Ramos (2003).
Let $T$ be a recipe tree with nodes $E_{T}$, where each node represents an element (either an ingredient or a cooking action). Let $N_{e}$ be the number of recipes in the repository that include element $e$. For each $e^{\prime}\in E_{T}\setminus\{e\}$, let $df_{e}(e^{\prime})$ denote the number of recipes containing both $e,e^{\prime}$. We define: [^3] To prevent score inflation due to typos or extremely rare ingredients, we exclude any ingredient that appears exceptionally infrequently in the entire dataset.
$idf_{e}(e’)=log\Bigl(\frac{N_{e}}{df_{e}(e^{\prime})}\Bigr)$
[…]
To compute novelty of an element $e$ in $T$, we compute its $idf_{e}$ score with respect to each of the other elements in $T$, and sum the top 10 values. To compute the overall novelty of $T$, we use the sum of the top-10 element-level novelty scores. A chicken and mango delight recipe (Figure 2) was deemed novel due to ingredient combinations like mango, crescent roll, nuts and broccoli, and actions such as steaming, blanching and unrolling.
@@ 4.5 Tree to Text (Step V)
After identifying the highest-ranked trees, we convert them back into natural-language recipes using an LLM. Similar to Section 4.2, we employ a chain-of-thought approach, instructing GPT-4o to translate the tree (encoded in DOT format) into a structured recipe, including a title, ingredient list, and step-by-step instructions. We then prompt the model to refine and correct the text for coherence, fluency, and consistency (see Appendix F for full prompts). Processing 1,000 recipe trees in this step required a total cost of approximately \$42.
We note that the LLM plays a key role in surface realization. Specifically, it draws on commonsense and domain knowledge to (1) fill in missing details (e.g., suggesting plausible ingredient quantities or cooking times), and (2) correct inconsistencies or omissions introduced during recombination (e.g., restoring a step for cooking raw chicken if it was omitted).
@@ 6 Evaluation
To answer this, we examine how our outputs compare to those of GPT-4o in two key aspects: diversity and creativity. For diversity, we explore whether our approach mitigates the well-documented issue of repetitiveness in LLMs. For creativity, we look for outputs that are both valuable (make sense) and novel (unexpected). We compare our model’s outputs with those of GPT-4o [^4] Version: gpt-4o-2024-08-06 (latest stable version of the model at the time of writing this paper). on two tasks:
Experiment 1. We evaluated GPT-4o and $\mathcal{D}$ishCover on their ability to generate creative recipes combining a given dish pair. The evaluation included 10 randomly selected dish pairs. For each pair, we selected 5 recipes generated by GPT-4o and compared them to the top 5 recipes generated by $\mathcal{D}$ishCover (50 recipes per model).
Experiment 2. To broaden the scope of our analysis, we wish to evaluate the most creative recipes GPT-4o and $\mathcal{D}$ishCover could generate in general, without limiting them to a given input pair. We used 100 recipes from each model. For $\mathcal{D}$ishCover, we selected 100 recipes from our 5K-recipe dataset, which includes a novelty score for each generated recipe. To ensure that the dishes do not come from a very small number of inputs, we employed simulated annealing Bertsimas and Tsitsiklis (1993), maximizing recipe novelty while enforcing a constraint about the maximum number of appearances of each dish.